Solve the problem involving the third law of Kepler's planetary motion with complete solutions. Mars takes 1.88 years to complete one orbit around the sun. What is its average distance from the sun?

1
Expert's answer
2021-12-20T17:57:24-0500

Let's apply the Third Kepler's Law of planetary motion:


"\dfrac{P_M^2}{a_M^3}=\dfrac{P_E^2}{a_E^3},""a_M=\sqrt{\dfrac{P_M^2a_E^3}{P_E^2}},""a_M=\sqrt[3]{\dfrac{(1.88\ yr)^2\times(1.5\times10^{11}\ m)^3}{(1\ yr)^2}}=2.285\times10^8\ km."